# Bayesian Inferential Motion Planning Using Heavy-Tailed Distributions

### v1

@@ I Introduction
Bayesian inferential motion planning (BIMP) has recently attracted a growing interest due to its promise to address some complex problems. BIMP focuses on inferring the best motion plans using the planning objectives and constraints as the evidence within a Bayesian framework berntorp2019motion ; askari2023model . It establishes a prior probability for motion plans, considers the likelihood of achieving the planning objectives and satisfying constraints given these plans, and then applies Bayes’ rule to obtain the posterior probability of the plans. BIMP harnesses the power of probabilistic inference to offer multi-fold advantages. It allows to easily incorporate robot models into the predictive planning process and enables the use of a plentiful number of nonlinear state estimation for motion planning. Especially, Monte Carlo sampling-based inference has gained significant use. For example, the studies in berntorp2017path ; askari2024motion exploit particle filtering/smoothing and ensemble Kalman smoothing to estimate the trajectories for autonomous vehicles. What makes sampling-based BIMP even more appealing is its capability of tackling complicated problems. For example, motion planning with machine learning dynamic models is challenging due to its complex optimization landscapes, but well-designed sampling schemes can achieve remarkably high computational efficiency askari2023model ; askari2024motion . Further, the idea of Bayesian inference has shown effectiveness in dealing with various tricky problems in motion planning, e.g., probabilistic collision avoidance and uncertain or incomplete knowledge of the environment patil2012estimating .
A key question for sampling-based BIMP is how to achieve high-efficiency search in probabilistic state spaces for high-quality motion plans. To address this question, the literature has introduced new sampling-based inference methods. In askari2023model , implicit particle filtering/smoothing seeks to concentrate sampling-based search within high-probability regions to improve both sampling efficiency and planning accuracy; in askari2024motion , a sequential single-pass ensemble Kalman smoothing is used to speed up computation. For sampling-based search, a main challenge is that high-quality motion plans may lie in the less probable regions of the prior probability. Particle filtering/smoothing can partly handle the situation, yet demanding the use of many particles at more computational expenses. The challenge is realistically impactful for ensemble Kalman smoothing, as its dependence on Gaussian distribution assumptions would hold it back from searching broadly.
@@ II BIMP Problem Setup
where $\bm{x}$ is the state including the position, heading angle, and velocity, and $\bm{u}$ is the input including the acceleration and steering angle. In this paper, we specifically use a neural network dynamic model for $f(\cdot,\cdot)$ as in askari2023model , leveraging its data-driven learning capability to grasp vehicle dynamics. This vehicle must move in such a way that $\bm{x}$ tracks a reference $\bm{r}$, generated by a higher-level decision-making module, while its input $\bm{u}$ follows $\bm{s}$. Some other so-called obstacle vehicles (OVs) are driving on the road, with their state denoted as $\bm{x}_{k}^{\mathrm{OV},i}$ for $i=1,2,\ldots,N_{O}$, where $N_{O}$ is the number of OVs. The EV must avoid colliding with the OVs, thus imposing the following constraint in motion planning:
@@
In above, $\bm{x}$, $\bm{u}$ and $\Delta\bm{u}$ together represent the state of this auxiliary system. The virtual observations of $\bm{x}$ and $\bm{u}$ are $\bm{r}$ and $\bm{s}$, due to the anticipation that the vehicle will the references. Further, all the constraints in (2)-(5) are summarized as
$\bm{\phi}(\bm{x},\bm{u},\bm{\Delta}\bm{u})\leq\bm{0}.$ (8)
@@
$\bar{\bm{x}}_{k}=\begin{bmatrix}{\bm{x}}_{k}\cr{\bm{u}}_{k}\cr{\bm{\Delta}\bm{u}}_{k}\cr\end{bmatrix},\ \bar{\bm{y}}_{k}=\begin{bmatrix}\bm{r}_{k}\cr\bm{s}_{k}\cr\bm{z}_{k}\end{bmatrix},\ \bar{\bm{w}}_{k}=\begin{bmatrix}\bm{0}\cr{\bm{w}}_{k}\cr{\bm{w}}_{k}\cr\end{bmatrix},\ \bar{\bm{v}}_{t}=\begin{bmatrix}\bm{v}_{\bm{x},k}\cr\bm{v}_{\bm{u},k}\cr\bm{v}_{\bm{z},k}\end{bmatrix},$
@@
By design, this framework attempts to infer the motion plans that have the maximum posterior probability conditioned on the driving objectives and constraints. What is shown in is a state smoothing problem, which is has analytical solution for the nonlinear $\bar{f}(\cdot)$. In the literature, this problem has been addressed using particle filtering/smoothing and ensemble Kalman smoothing askari2024motion ; berntorp2019motion . Next, we will develop a new method that builds upon heavy-tailed distributions.
@@ III BIMP Using Student’s-$t$ Distributions
where $\bar{\bm{w}}^{j}_{t-1}$ are drawn from $St\left(\bm{0};\bm{\Sigma}_{\bm{w}},\nu_{\bm{w}}\right)$. We join $\bar{\bm{x}}^{j}_{t|t-1}$ to $\bm{\mathcal{X}}^{j}_{t-1|t-1}$ via the concatenating operation:
@@ IV Connection to Nonlinear Model Predictive Control
$+\sum_{t=k}^{k+H}(-\frac{\nu_{\bm{\Delta}\bm{u}}+n_{\Delta\bm{u}}}{2})\log\Big{[}1+\frac{1}{\nu_{\Delta\bm{u}}}\Delta\bm{u}_{t}^{\top}\bm{\Sigma}_{\bm{w}}^{-1}\Delta\bm{u}_{t}\Big{]}.$

### v2

@@ I Introduction
Bayesian inferential motion planning (BIMP) has recently attracted a growing interest due to its promise to address some complex problems. BIMP focuses on inferring the best motion plans using the planning objectives and constraints as the evidence within a Bayesian framework berntorp2019motion ; askari2025model . It establishes a prior probability for motion plans, considers the likelihood of achieving the planning objectives and satisfying constraints given these plans, and then applies Bayes’ rule to obtain the posterior probability of the plans. BIMP harnesses the power of probabilistic inference to offer multi-fold advantages. It allows to easily incorporate robot models into the predictive planning process and enables the use of a plentiful number of nonlinear state estimation for motion planning. Especially, Monte Carlo sampling-based inference has gained significant use. For example, the studies in berntorp2017path ; askari2024motion exploit particle filtering/smoothing and ensemble Kalman smoothing to estimate the trajectories for autonomous vehicles. What makes sampling-based BIMP even more appealing is its capability of tackling complicated problems. For example, motion planning with machine learning dynamic models is challenging due to its complex optimization landscapes, but well-designed sampling schemes can achieve remarkably high computational efficiency askari2025model ; askari2024motion . Further, the idea of Bayesian inference has shown effectiveness in dealing with various tricky problems in motion planning, e.g., probabilistic collision avoidance and uncertain or incomplete knowledge of the environment patil2012estimating .
A key question for sampling-based BIMP is how to achieve high-efficiency search in probabilistic state spaces for high-quality motion plans. To address this question, the literature has introduced new sampling-based inference methods. In askari2025model , implicit particle filtering/smoothing seeks to concentrate sampling-based search within high-probability regions to improve both sampling efficiency and planning accuracy; in askari2024motion , a sequential single-pass ensemble Kalman smoothing is used to speed up computation. For sampling-based search, a main challenge is that high-quality motion plans may lie in the less probable regions of the prior probability. Particle filtering/smoothing can partly handle the situation, yet demanding the use of many particles at more computational expenses. The challenge is realistically impactful for ensemble Kalman smoothing, as its dependence on Gaussian distribution assumptions would hold it back from searching broadly.
@@ II BIMP Problem Setup
where $\bm{x}$ is the state including the position, heading angle, and velocity, and $\bm{u}$ is the input including the acceleration and steering angle. In this paper, we specifically use a neural network dynamic model for $f(\cdot,\cdot)$ as in askari2025model , leveraging its data-driven learning capability to grasp vehicle dynamics. This vehicle must move in such a way that $\bm{x}$ tracks a reference $\bm{r}$, generated by a higher-level decision-making module, while its input $\bm{u}$ follows $\bm{s}$. Some other so-called obstacle vehicles (OVs) are driving on the road, with their state denoted as $\bm{x}_{k}^{\mathrm{OV},i}$ for $i=1,2,\ldots,N_{O}$, where $N_{O}$ is the number of OVs. The EV must avoid colliding with the OVs, thus imposing the following constraint in motion planning:
@@
In above, $\bm{x}$, $\bm{u}$ and $\Delta\bm{u}$ together represent the state of this auxiliary system. The virtual observations of $\bm{x}$ and $\bm{u}$ are $\bm{r}$ and $\bm{s}$, due to the anticipation that the vehicle will track the references. Further, all the constraints in (2)-(5) are summarized as
$\bm{\phi}(\bm{x},\bm{u},\Delta\bm{u})\leq\bm{0}.$ (8)
@@
$\bar{\bm{x}}_{k}=\begin{bmatrix}{\bm{x}}_{k}\cr{\bm{u}}_{k}\cr{\Delta\bm{u}}_{k}\cr\end{bmatrix},\ \bar{\bm{y}}_{k}=\begin{bmatrix}\bm{r}_{k}\cr\bm{s}_{k}\cr\bm{z}_{k}\end{bmatrix},\ \bar{\bm{w}}_{k}=\begin{bmatrix}\bm{0}\cr{\bm{w}}_{k}\cr{\bm{w}}_{k}\cr\end{bmatrix},\ \bar{\bm{v}}_{t}=\begin{bmatrix}\bm{v}_{\bm{x},k}\cr\bm{v}_{\bm{u},k}\cr\bm{v}_{\bm{z},k}\end{bmatrix},$
@@
By design, this framework attempts to infer the motion plans that have the maximum posterior probability conditioned on the driving objectives and constraints. What is shown in (10) is a state smoothing problem, which is has analytical solution for the nonlinear $\bar{f}(\cdot)$. In the literature, this problem has been addressed using particle filtering/smoothing and ensemble Kalman smoothing askari2024motion ; berntorp2019motion . Next, we will develop a new method that builds upon heavy-tailed distributions.
@@ III BIMP Using Student’s-$t$ Distributions
where $\bar{\bm{w}}^{j}_{t-1}$ are drawn from $St\left(\bm{0};\bm{\Sigma}_{\bm{w}},\nu_{\bm{w}}\right)$. We join $\bar{\bm{x}}^{j}_{t|t-1}$ to $\bm{\mathcal{X}}^{j}_{t-1|t-1}$ via the concatenating operation
@@ IV Connection to Nonlinear Model Predictive Control
$+\sum_{t=k}^{k+H}(-\frac{\nu_{\Delta\bm{u}}+n_{\Delta\bm{u}}}{2})\log\Big{[}1+\frac{1}{\nu_{\Delta\bm{u}}}\Delta\bm{u}_{t}^{\top}\bm{\Sigma}_{\bm{w}}^{-1}\Delta\bm{u}_{t}\Big{]}.$